OpenStudy (anonymous):

A 90.1-kg baseball player slides into second base. The coefficient of kinetic friction between the player and the ground is μk = 0.511. (a) What is the magnitude of the frictional force? (b) If the player comes to rest after 1.91 s, what is his initial speed?

OpenStudy (anonymous):

The frictional force is simply\[F_f = \mu_k m g\] His initial speed can be determined from Impulse\[I = \int\limits F dt\]From Newton's Second Law\[F = {dp \over dt}\]where p is momentum. Therefore,\[I = \int\limits {dp \over dt} dt = \int\limits dp = \Delta p = m(v_f - v_i)\]where \(v_f\) is the final velocity and \(v_i\) is the initial velocity. We can find the initial velocity of the player as\[F \Delta t = mv_i \rightarrow \mu_k m g \Delta t = m v_i \rightarrow \mu_k g \Delta t = v_i\]